OpenStudy (anonymous):

Factor the following Y^4-16Y^3 Can someone explain why it is y = 0 and =16 The way i factored it i got y^2 (y-4)(y+4) So i would get y = 0 y= 4 y=-4 Why is this wrong

OpenStudy (saifoo.khan):

y^3(y-16)

OpenStudy (saifoo.khan):

y = 0 & y = 16

OpenStudy (anonymous):

my answer yields the same results so why is one wrong and one not? Thanks by the way

OpenStudy (saifoo.khan):

that is wrong b/c y^3 is the common. so you are left with y-16 so: Y^3(y-16) = 0 y^3 = 0 y = 0 y-16 = 0 y = 16

OpenStudy (saifoo.khan):

are you getting it now?

OpenStudy (anonymous):

Thanks

OpenStudy (saifoo.khan):

Welks.